# Multi-Agent Target Assignment and Path Finding for Intelligent Warehouse: A Cooperative Multi-Agent Deep Reinforcement Learning Perspective

### v2

@@ Abstract
Multi-agent target assignment and path planning (TAPF) are two key problems in intelligent warehouses. However, most literature only addresses one of these two problems separately. In this study, we propose a method to simultaneously solve target assignment and path planning from the perspective of cooperative multi-agent deep reinforcement learning (RL). This paper models the TAPF problem for intelligent warehouses to cooperative multi-agent deep RL and simultaneously addresses TAPF based on multi-agent deep RL. Furthermore, previous literature rarely considers the physical dynamics of agents. In this study, the physical dynamics of the agents are considered. Experimental results show that our method performs well in various task settings, which means that the target assignment is solved reasonably well and the planned path is almost the shortest. Moreover, our method is more time-efficient than baselines.
keywords:
Intelligent warehouse, Task assignment, Path finding, Multi-agent reinforcement learning
journal: Applied Intelligence
1 Introduction
@@
With the rapid development of logistics-related industries, new opportunities and challenges have been proposed for intelligent warehousing [1, 2]. Traditional warehouse technologies use conveyor belts and other equipment to complete material handling. They are inflexible and not easily extensible. Intelligent warehousing utilizes a multi-agent system to deliver goods to designated locations, greatly improving warehousing efficiency. The intelligent warehouse system is divided mainly into the order fulfillment system, such as the kiva system [3] and the logistics sorting center [4]. In order fulfillment systems, mobile agents move inventory pods to inventory stations. Then, workers take goods from inventory pods, and agents move the inventory pods to their original locations. In logistics sorting centers, agents take goods from loading stations and deliver them to chutes in the center of warehouse. The approach proposed in this study is based on the logistics sorting center.
@@
Task assignment and path finding (TAPF) are two important processes in intelligent warehousing. The system first assigns a specified task to each agent according to the order requirements. Then the agent transports the goods from the origin to the destination and ensures that the path does not conflict with other agents. The task assignment and path planning problem is typically NP-hard [5], which has a large search space. Thus, directly solving this problem is difficult. In general, two steps are required to solve the TAPF problem. The first step is multi-agent task assignment (MATA), assigning tasks to agents without considering possible path conflicts between agents. The second step is path planning for all agents by using multi-agent path finding (MAPF) algorithms. Numerous studies in the literature study MATA and MAPF; we describe these methods in detail in the next section. Although solving TAPF separately reduces the difficulty of the total problem, this type of method ignores the mutual influence between task assignment and path planning. Reasonable task assignment can not only effectively reduce the path length of warehouse agents and improve the operation efficiency, but also help avoid path conflicts between different agents. Therefore, it is necessary to solve the TAPF problems together. In the TAPF problem, we assume that each idle agent and inventory pod are homogeneous. Thus, we can assign any task to any agent.
@@
Recently, deep reinforcement learning (RL) has received great attention due to its remarkable successes in widely challenging domains, such as robotic control [6], multi-agent problems [7], and video games [8]. In this study, we introduce cooperative multi-agent deep RL to solve the TAPF problem to address the TAPF problem simultaneously.
The main contributions of this study can be summarized as follows:
  • 1.
@@
    This paper models the TAPF problem for intelligent warehouses to cooperative multi-agent deep RL and simultaneously addresses TAPF based on multi-agent deep RL
  • 2.
@@
    In this paper, the physical dynamics of the agents are considered in the path planning phase.
  • 3.
@@
    Experimental results show that our method performs well in various of task settings, which means the target assignment is solved reasonably and the planned path is almost the shortest. Furthermore, our method is more time-efficient than the baselines.
The rest of this study is organized as follows. Section 2 discusses the related work of TAPF in an intelligent warehouse. Section 3 introduces the preliminaries of the Markov Decision Process, RL, the cooperative multi-agent deep RL, and the definition of MAPF and TAPF problem. Section 4 detailedly describes our method. Section 5 provides experimental results to verify the improved efficiency of our method. Section 6 presents conclusions and future work.
2 Related Work
2.1 Multi-Agent Task Assignment
@@
Through multi-agent task assignment algorithms, we can maximize the utilization of warehouse resources [9]. At present, multi-agent task assignment algorithms in intelligent warehouses can be classed into centralized and distributed classes according to the management mode. In the centralized task assignment class, a central control system is set up, which is responsible for task assignment, assigning tasks assigned to agents for execution [10]. Classical centralized task assignment algorithms include: Hungarian algorithm [11], tabu search algorithm [12], genetic algorithm [10], etc.. In the distributed assignment class, each agent in the warehouse plans its own task sequence according to tasks and environmental information [13]. This class method effectively reduces the load of the central control system and is more flexible and adaptable, but it may not find the global optimal solution [14]. Furthermore, the distributed assignment class mainly includes: learning-based methods [15] and market auction methods [16].
2.2 Multi-Agent Path Finding
@@
In recent years, MAPF has become a hot research direction in the fields of computer science and robotics. Classical MAPF algorithms can be divided into optimal and sub-optimal types according to whether the solution results meet the optimality. The optimal typical MAPF algorithms include A*-based [17], conflict-based search [18], increasing cost tree search based [19], and compilation-based [20] methods. Sub-optimal classical MAPF algorithms include search-based [21], rule-based [22], compilation-based [23], bounded conflict-based search [24], and genetic algorithm-based [25] methods. However, traditional MAPF algorithms have poor real-time performance. Therefore, numerous researchers are beginning to study MAPF based on deep RL [26] to solve this problem. PRIMAL [27] is a typical MAPF algorithm based on deep RL. However, PRIMAL still solves the multi-agent problem by single-agent deep RL method [27]. In this study, we propose a method to simultaneously address target assignment and path planning from the perspective of cooperative multi-agent deep RL. Furthermore, in the simulation environment, the above-mentioned various MAPF algorithms rarely consider the physical dynamics of agents. In this study, the physical dynamics factors of the agent are considered.
2.3 Cooperative Multi-Agent Deep Reinforcement Learning
@@
Cooperative multi-agent RL deals with systems consisting of several agents which interact in a common environment. Cooperative multi-agent RL aims to learn a policy that can make all agents work together to complete a common goal. Multi-agent deep RL has received great attention due to its capability to allow agents to learn to make decisions in multi-agent environments through interactions. Independent Q-learning (IQL) [28] solves the multi-agent RL problem by decomposing it into a collection of simultaneous single-agent RL problems that share the same environment. However, IQL cannot solve the nonstationarity [29] problem caused by the changing policies of other agents. Value-Decomposition Networks (VDN) [30] proposes to learn to linearly decompose a team value function into a series of per-agent value functions. After that, QMIX [7] utilizes a mixing network architecture to approximate all agents’ joint state-action value. The mixing network non-linearly combines each agent’s state-action value, which is only a condition on each agent’s local observations. Furthermore, QMIX enforces the joint state-action value monotonic with each agent state-action value by employing a network structural monotonicity constraint to the mixing network.
@@
However, these above algorithms cannot handle continuous action space problems. The deep deterministic policy gradient (DDPG) [6] is a representative method in the continuous action task for a single agent. Multi-agent deep deterministic policy gradient (MADDPG) [31] extended DDPG from single agent setting to the multi-agent setting by using centralized training and decentralized execution paradigm [32]. MADDPG is a representative algorithm for multi-agent continuous action problems. Thus, in this study, we model the TAPF in the intelligent warehouse as a cooperative multi-agent deep RL problem and introduce MADDPG to address this problem simultaneously.
3 Preliminaries
This section provides a summary of Markov Decision Process, RL, cooperative multi-agent deep RL, and the definition of MAPF and TAPF problems.
3.1 Markov Decision Process and RL
@@
This study considers a finite-horizon Markov Decision Process [26], defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,T)$. $\mathcal{S}$ denotes the state space, $\mathcal{A}$ represents the finite action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the state transition distribution, $r:\mathcal{S}\times\mathcal{A}\rightarrow R$ denotes the reward function, $\gamma\in[0,1)$ denotes the discount factor and $T$ is a time horizon. At each time step $t$, an action $a_{t}\in A$ is chosen from a policy $\pi$. After transiting into the next state by sampling from $\mathcal{P}\left(s_{t+1}\mid s_{t},a_{t}\right)$, the agent obtains a immediate reward $r\left(s_{t},a_{t}\right)$. The agent continues performing actions until it enters a terminal state or $t$ reaches the time horizon $T$. RL aims to learn the policy $\pi:S\times A\rightarrow[0,1]$ for decision-making problems by maximizing discounted cumulative rewards $\mathbb{E}_{\pi,\mathcal{P}}\left[r_{0:T-1}\right]=\mathbb{E}_{\pi,\mathcal{P}}\left[\sum_{0}^{T-1}\gamma^{t}r\left(s_{t},a_{t}\right)\right]$. In this study, we do not have access to the environment dynamics $\mathcal{P}$, which means model-free deep RL learning.
3.2 Cooperative Multi-Agent Deep RL
@@
A fully cooperative multi-agent problem can be modeled as a Markov Decision Process [26], described by a tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},N,\gamma,T)$. $s\in S$ represents the true state. At each timestep $t$, each agent $n\in N:=\{1,\ldots,N\}$ selects an action $a^{n}_{t}\in\mathcal{A}$, then the joint action $\boldsymbol{a_{t}}=\{a^{1}_{t},a^{2}_{t},...,a^{N}_{t}\}$ is obtained. $\mathcal{P}\left(s^{\prime}_{t+1}\mid s_{t},\boldsymbol{a_{t}}\right):\mathcal{S}\times\mathcal{A}^{N}\times S\rightarrow[0,1]$ denotes the state transition function. $r(s_{t},\boldsymbol{a_{t}}):\mathcal{S}\times\mathcal{A}^{N}\rightarrow R$ denotes the reward function shared by all agents, $\gamma\in[0,1)$ represents the discount factor, and $T$ denotes the time horizon. Each agent $n$ has an independent observation $o\in\mathcal{O}$ and a history of action observation $\tau^{n}\in\mathcal{T}\equiv(\mathcal{O}\times\mathcal{A})$. A stochastic policy $\pi^{n}\left(a^{n}\mid\tau^{n}\right):\mathcal{T}\times A\rightarrow[0,1]$ is based on the history of action-observation. Let $\pi^{n}$ denote agent $n$’ policy; cooperative agents aim to maximize:
@@
(a) five agents - five tasks
(b) five agents - twenty tasks
@@
3.3 Definition of MAPF and TAPF problem
MAPF problem: In the multi-agent path finding (MAPF) problem, the input includes an environment consisting of N agents. The environment defines all possible start and goal positions for the agents. Each agent $n$ has its own start position $start^{n}$ and goal position $goal^{n}$. In MAPF, time is typically assumed to be discrete. At time $t$, agent $n$ is at a specific location and can execute an action $a^{n}_{t}$, subsequently moving to a new position or staying at the current position. The action can be represented as $a^{n}_{t}:\text{current position}\rightarrow\text{new position}$. The path $\pi_{n}$ of agent $n$ consists of a sequence of actions $\pi_{n}=(a^{n}_{1},a^{n}_{2},\ldots,a^{n}_{T-1})$. Conflicts in MAPF typically involve position and movement collisions between agents.
@@ 3.3 Definition of MAPF and TAPF problem
4 Target Assignment and Path Finding for Intelligent Warehouse: A Cooperative Multi-Agent Deep RL Perspective
In Section 4.1, we detailedly describe our method that models the TAPF problem as a cooperative multi-agent RL problem. In Section 4.2, we introduce the MADDPG algorithm to solve TAPF.
4.1 Modeling TAPF as a MARL Problem
As shown in Figure 1, we provide two scenarios to describe how we model the TAPF problem (described in Section 3.3) as a cooperative MARL problem. As introduced in Section 3.2, cooperative multi-agent deep RL can be modeled as a tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},N,\gamma,T)$. In this study, the state transition function $\mathcal{P}\left(s^{\prime}_{t+1}\mid s_{t},\boldsymbol{a_{t}}\right):\mathcal{S}\times\mathcal{A}^{N}\times\mathcal{S}\rightarrow[0,1]$ is unknown, which means the agent cannot get the environment dynamics. This is consistent with the real-world TAPF problem. The number of agents $N$ in this study can be set to any integer number. The discount factor $\gamma=0.99$ is the most common setting in deep RL. Since the main elements in multiagent deep RL are the observation space $\mathcal{O}$ (or state space $\mathcal{S}$), the action space $\mathcal{A}$, and the reward function $r$. Thus, we describe these elements in detail.
Figure 2: The physical dynamics of agents

### v3

@@ Abstract
Multi-agent target assignment and path planning (TAPF) are two key problems in intelligent warehouses. However, most literature only addresses one of these two problems separately. This paper proposes a method to simultaneously solve target assignment and path planning from the cooperative multi-agent deep reinforcement learning (RL) perspective. This paper models the TAPF problem for intelligent warehouses to cooperative multi-agent deep RL and simultaneously addresses task assignment and path finding based on multi-agent deep RL. Furthermore, previous literature rarely considers the physical dynamics of agents. In this paper, the physical dynamics of the agents are considered. This is more in line with actual robot movement. Experimental results show that the proposed method performs well in various task settings, which means that the target assignment is solved reasonably well and the planned path is almost the shortest. Moreover, the proposed method is more time-efficient than baselines.
I Introduction
@@
With the rapid development of logistics-related industries, new opportunities, and challenges have been proposed for intelligent warehousing [1, 2]. Traditional warehouse technologies use conveyor belts and other equipment to complete material handling. They are inflexible and not easily extensible. Intelligent warehousing utilizes a multi-agent system to deliver goods to designated locations, greatly improving efficiency. The intelligent warehouse system is divided mainly into the order fulfillment system, such as the kiva system [3] and the logistics sorting center [4]. In order fulfillment systems, mobile agents move inventory pods to inventory stations. Then, workers take goods from inventory pods, and agents move the inventory pods to their original locations. In logistics sorting centers, agents take goods from loading stations and deliver them to chutes in the center of the warehouse. The approach proposed in this study is based on the logistics sorting center.
@@
Task assignment and path finding (TAPF) are two important processes in intelligent warehousing. The system first assigns a specified task to each agent according to the order requirements. Then, the agent transports the goods from the origin to the destination and ensures that the path does not conflict with other agents. The task assignment and path planning problem is typically NP-hard [5], which has a large search space. Thus, directly solving this problem is difficult. In general, two steps are required to solve the TAPF problem. The first step is multi-agent task assignment (MATA), assigning tasks to agents without considering possible path conflicts between agents. The second step is path planning for all agents by using multi-agent path finding (MAPF) algorithms. Numerous studies in the literature study MATA and MAPF; we describe these methods in detail in the next section. Although solving TAPF separately reduces the difficulty of the total problem, this type of method ignores the mutual influence between task assignment and path planning. Reasonable task assignments can reduce warehouse agents’ path length, improve operation efficiency, and help avoid path conflicts between agents. Therefore, it is necessary to solve the TAPF problems together. In the TAPF problem, we assume that each idle agent and inventory pod are homogeneous. Thus, we can assign any task to any agent.
@@
Recently, deep reinforcement learning (RL) has received great attention due to its remarkable successes in widely challenging domains, such as robotic control [6, 7, 8], multi-agent problems [9], and video games [10]. In this study, we introduce cooperative multi-agent deep RL to solve the TAPF problem and address it simultaneously.
The main contributions of this paper can be summarized as follows:
@@
  • This paper proposes a method to simultaneously solve target assignment and path planning from the cooperative multi-agent deep RL perspective. This paper models the TAPF problem for intelligent warehouses to cooperative multi-agent deep RL and simultaneously solves task assignment and path finding based on the multi-agent deep RL algorithm.
@@
  • Furthermore, in this paper, the physical dynamics of the agents are considered in the path planning phase. This is more in line with actual robot movement.
@@
  • Experimental results show that the proposed method performs well in various task settings. This means the target assignment is solved reasonably, and the planned path is almost the shortest. Furthermore, the proposed method is more time-efficient than the baselines.
II Related Work
II-A Multi-Agent Task Assignment
@@
Through multi-agent task assignment algorithms, we can maximize the utilization of warehouse resources [11]. Currently, multi-agent task assignment algorithms in intelligent warehouses can be classed into centralized and distributed classes according to the management mode. In the centralized task assignment class, a central control system is set up, which is responsible for task assignment, assigning tasks assigned to agents for execution [12]. Classical centralized task assignment algorithms include: Hungarian algorithm [13], tabu search algorithm [14], genetic algorithm [12], etc. In the distributed assignment class, each agent in the warehouse plans its own task sequence according to tasks and environmental information [15]. This class method effectively reduces the load of the central control system and is more flexible and adaptable, but it may not find the global optimal solution [16]. Furthermore, the distributed assignment class mainly includes learning-based methods [17] and market auction methods [18].
II-B Multi-Agent Path Finding
@@
In recent years, MAPF has become a hot research direction in computer science and robotics. Classical MAPF algorithms can be divided into optimal and sub-optimal types according to whether the solution results meet the optimality. The optimal typical MAPF algorithms include A*-based [19], conflict-based search [20], increasing cost tree search based [21], and compilation-based [22] methods. Sub-optimal classical MAPF algorithms include search-based [23], rule-based [24], compilation-based [25], bounded conflict-based search [26], and genetic algorithm-based [27] methods. However, traditional MAPF algorithms have poor real-time performance. Therefore, numerous researchers are beginning to study MAPF based on deep RL [28] to solve this problem. PRIMAL [29] is a typical MAPF algorithm based on deep RL. However, PRIMAL still solves the multi-agent problem by single-agent deep RL method [29]. In this study, we propose a method to simultaneously address target assignment and path planning from the cooperative multi-agent deep RL perspective. Furthermore, in the simulation environment, the above-mentioned various MAPF algorithms rarely consider the physical dynamics of agents. In this study, the physical dynamics factors of the agent are considered.
II-C Cooperative Multi-Agent Deep RL
@@
Cooperative multi-agent RL (MARL) deals with systems consisting of several agents that interact in a common environment. Cooperative multi-agent RL aims to learn a policy to make all agents work together to complete a common goal. Multi-agent deep RL has received great attention due to its capability to allow agents to learn to make decisions in multi-agent environments through interactions. Independent Q-learning (IQL) [30] solves the multi-agent RL problem by decomposing it into a collection of simultaneous single-agent RL problems that share the same environment. However, IQL cannot solve the nonstationarity [31] problem caused by the changing policies of other agents. Value-Decomposition Networks (VDN) [32] proposes to learn to linearly decompose a team value function into a series of per-agent value functions. After that, QMIX [9] utilizes a mixing network architecture to approximate all agents’ joint state-action value. The mixing network non-linearly combines each agent’s state-action value, which is only a condition on each agent’s local observations. Furthermore, QMIX enforces the monotonic joint state-action value with each agent state-action value by employing a network structural monotonicity constraint to the mixing network.
@@
However, these algorithms cannot handle continuous action space problems. The deep deterministic policy gradient (DDPG) [6] is a representative method in the continuous action task for a single agent. Multi-agent deep deterministic policy gradient (MADDPG) [33] extended DDPG from single agent setting to the multi-agent setting by using centralized training and decentralized execution paradigm [34]. MADDPG is a representative algorithm for multi-agent continuous action problems. Thus, in this study, we model the TAPF in the intelligent warehouse as a cooperative multi-agent deep RL problem and use MADDPG to solve this problem simultaneously.
III Preliminaries
III-A Markov Decision Process and RL
@@
This study considers a finite-horizon Markov decision process [28], defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,T)$. $\mathcal{S}$ denotes the state space, $\mathcal{A}$ represents the finite action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the state transition distribution, $r:\mathcal{S}\times\mathcal{A}\rightarrow R$ denotes the reward function, $\gamma\in[0,1)$ denotes the discount factor and $T$ is a time horizon. At each time step $t$, an action $a_{t}\in A$ is chosen from a policy $\pi$. After transiting into the next state by sampling from $\mathcal{P}\left(s_{t+1}\mid s_{t},a_{t}\right)$, the agent obtains a immediate reward $r\left(s_{t},a_{t}\right)$. The agent continues performing actions until it enters a terminal state or $t$ reaches the time horizon $T$. RL aims to learn the policy $\pi:S\times A\rightarrow[0,1]$ for decision-making problems by maximizing discounted cumulative rewards $\mathbb{E}_{\pi,\mathcal{P}}\left[r_{0:T-1}\right]=\mathbb{E}_{\pi,\mathcal{P}}\left[\sum_{0}^{T-1}\gamma^{t}r\left(s_{t},a_{t}\right)\right]$. In this study, we cannot access the environment dynamics $\mathcal{P}$, which means model-free deep RL learning.
III-B Cooperative Multi-Agent Deep RL
@@
A fully cooperative multi-agent problem can be modeled as a Markov decision process [28], described by a tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},N,\gamma,T)$. $s\in S$ represents the true state. At each timestep $t$, each agent $n\in N:=\{1,\ldots,N\}$ selects an action $a^{n}_{t}\in\mathcal{A}$, then the joint action $\bm{a_{t}}=\{a^{1}_{t},a^{2}_{t},...,a^{N}_{t}\}$ is obtained. $\mathcal{P}\left(s^{\prime}_{t+1}\mid s_{t},\bm{a_{t}}\right):\mathcal{S}\times\mathcal{A}^{N}\times S\rightarrow[0,1]$ denotes the state transition function. $r(s_{t},\bm{a_{t}}):\mathcal{S}\times\mathcal{A}^{N}\rightarrow R$ denotes the reward function shared by all agents, $\gamma\in[0,1)$ represents the discount factor, and $T$ denotes the time horizon. Each agent $n$ has an independent observation $o\in\mathcal{O}$ and a history of action observation $\tau^{n}\in\mathcal{T}\equiv(\mathcal{O}\times\mathcal{A})$. A stochastic policy $\pi^{n}\left(a^{n}\mid\tau^{n}\right):\mathcal{T}\times A\rightarrow[0,1]$ is based on the history of action-observation. Let $\pi^{n}$ denote agent $n$’ policy; cooperative agents aim to maximize:
@@
Figure 2: The physical dynamics of agents
III-C Definitions of MAPF and TAPF Problems
MAPF problem: In the multi-agent path finding (MAPF) problem, the input includes an environment consisting of N agents. The environment defines all possible start and goal positions for the agents. Each agent $n$ has its own start position $start^{n}$ and goal position $goal^{n}$. At time $t$, agent $n$ is at a specific location and can execute an action $a^{n}_{t}$, subsequently moving to a new position or staying at the current position. The action can be represented as $a^{n}_{t}:\text{current position}\rightarrow\text{new position}$. The path $\pi_{n}$ of agent $n$ consists of a sequence of actions $\pi_{n}=(a^{n}_{1},a^{n}_{2},\ldots,a^{n}_{T-1})$. Conflicts in MAPF typically involve position and movement collisions between agents.
@@ III-C Definitions of MAPF and TAPF Problems
IV Target Assignment and Path Finding for Intelligent Warehouse: A Cooperative Multi-Agent Deep RL Perspective
In Section IV-A, we detailedly describe our method that models the TAPF problem as a cooperative multi-agent RL problem. In Section IV-B, we introduce the MADDPG algorithm to solve TAPF.
IV-A Modeling TAPF as a MARL Problem
As shown in Fig. 1, we provide two scenarios to describe how we model the TAPF problem (described in Section III-C) as a cooperative MARL problem. As introduced in Section III-B, cooperative multi-agent deep RL can be modeled as a tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},N,\gamma,T)$. In this study, the state transition function $\mathcal{P}\left(s^{\prime}_{t+1}\mid s_{t},\bm{a_{t}}\right):\mathcal{S}\times\mathcal{A}^{N}\times\mathcal{S}\rightarrow[0,1]$ is unknown, which means the agent cannot get the environment dynamics. This is consistent with the real-world TAPF problem. The number of agents $N$ in this study can be set to any integer number. The discount factor $\gamma=0.99$ is the most common setting in deep RL. Since the main elements in multi-agent deep RL are the observation space $\mathcal{O}$ (or state space $\mathcal{S}$), the action space $\mathcal{A}$, and the reward function $r$. Thus, we describe these elements in detail.
Observation space: An agent’s observation contains the position and velocity itself, the relative positions of all tasks, the relative positions of the other agents, and the relative positions of neighboring obstacles. In this study, the number of other visible agents in an agent’s observation can be set to equal to or less than $N-1$. Taking Fig. 1 (a) as an example, the five purple circles represent five agents, and the five black circles represent five tasks. For obstacles, we set the agent to perceive the relative positions of the adjacent obstacles, and the number of adjacent obstacles can be set to any integer. In Fig. 1 (a) and (b), different color curves denote the navigation trajectories of different agents.